# A Causally Informed Pretraining Approach for Multimodal Foundation Models: Applications in Remote Sensing

### v3

@@ 4. Pretraining
As mentioned before, our proposed pretraining task is forecasting of satellite imagery using multimodal data in a causally informed fashion, a pretraining task that varies significantly from the traditional single modality reconstruction. Our proposed approach can be trained using the architecture described above in a direct shot, i.e initialise and update all layers at once. However, this may not be the best way to ensure that the $Emb_{STW}$ embeddings capture causal information, as that is the ultimate aim of our proposed approach. To ensure this information capture, we propose a phase wise pretraining process. Specifically we propose 2 phases of pretraining, where the first phase focuses on getting information from each modality and the second phase focuses on capturing the causal relationship between modalities in the embedding.
@@
To summarize, our proposed pretraining task is to predict a spectral image in the future (response) using a series of spectral images in the past (context) along with the weather till that future date (query). We call this pretraining task as Causally Informed Variable step Forecasting (CI-VSF), where the model is expected to forecast k steps into the future. To ensure that relationship across modalities is captured in the embeddings, we adopt a two phase pretraining process. Our hypothesis is that this extra knowledge infusion would help greatly in downstream tasks that rely on land growth and change dynamics such as crop prediction, land cover land use change, etc. A schematic of the stagewise pretraining can be found in appendix A.1. We mask out patches in the spectral imagery and weather data in all the stages to lead to better embeddings, which we describe in a future section(Sec 5.2).
5. Dataset
@@ 5.1. Data Sources
Our spectral imagery data comes from Sentinel imagery (Drusch et al., 2012) and our weather data is from ERA5 land analysis data (Hersbach et al., 2020). We chose these two sources due to their temporal resolutions and more importantly their availability globally from 2021. We randomly sampled around 10000 locations across land areas globally. Each location is of size 128x128 Sentinel pixels and for spectral imagery we collected all images for that region in that year. Due to missing data and improper coverage of some regions, the number of Sentinel samples from each region would vary. For example, regions in well covered regions such as US, might have upto 70 image instances in a year for that region, whereas regions like India (which is not as well covered) would have 40 image instances. For each instance, we collect six bands namely (B2, B3, B4, B8, B9, B12), which have shown to be the most useful in land cover related tasks and have been used in other works (Jakubik et al., 2023). For each image instance, we also collect the day of the year it came from, thus forming a series with values from 1 to 365 and a length the same as the number of image instances for that location. Though ERA5 data source consists of various bands that are useful for land cover related tasks we chose 5 bands that are the primary weather drivers, namely (temperature 2m min, temperature 2m max, total precipitation sum, $u$ component of wind 10m, $v$ component of wind 10m). ERA5 data is available at a daily temporal resolution and a spatial resolution of  11k meters.
To summarise, for each location our data comprises of:
$\bullet$ Spectral Imagery Series: A series of Sentinel2 Imagery each of 6 bands and of shape 128x128. Length of this series depends on coverage of the location.
$\bullet$ Weather Data Series: A series of ERA5 Land data of 5 bands and of shape 1x1, with a series length of 365 (one per day). The reason behind having only one pixel value from weather data is due to its coarse resolution (11km) compared to the area coverd by our Sentinel patch (1.28km x 1.28km)
@@ 5.2. Masking
Since we have a spatiotemporal input and architecture, we adopt a spatiotemporally uniform masking method, i.e, masking that is fair both spatially and temporally. In our unique masking strategy, there are an equal number of masked patches per timestamp as well as an equal number of masked patches per spatial patch location along the temporal axis. Figure 5 shows an example of such masking for a image series of 4 4x4 grids with 50% masking. From the Figure, we can see that in each timestamp image there are 8 patches masked and focusing on a particular patch location along the temporal dimension we notice that 2 patches are available. This ensures that all temporal patch series that can be created from unmasked patches at each spatial location would be of same length, easing the implementation of temporal components. Also the shared vision transformer will also have same number of outputs per timestamp due to same number of unmasked patches per timestamp.
@@ 6.1. Comparative Pretraining Frameworks
In our experiments, we evaluate the effectiveness of CI-VSF by comparing with several representative foundation modeling frameworks. We model these frameworks by varying the choices for input and pretraining tasks from existing works, as described below.
$\bullet$ SM-MR: Single Modality Masked Reconstruction (SM-MR) with satellite data as input and MR as the pretraining task. This is the most common setup used in existing remote sensing foundation models (e.g.,(Jakubik et al., 2023; Cong et al., 2022; Sun et al., 2022)).
$\bullet$ MM-MR: MultiModal Masked Reconstruction (MM-MR) with satellite and weather data as input series and MR as pretraining task. This represents the foundation models that use multiple modalities but still do masked reconstruction (eg. (Nedungadi et al., 2024; Tseng et al., 2023))
$\bullet$ SM-VSF: Single Modality Variable step Forecasting (SM-VSF) with the satellite data as input series and VSF as pretraining task.
It is important to note that the goal of the paper is not to release a new foundation model but show efficacy of our novel pre-training strategy for building foundation models. We are aiming to show how a causally guided pretraining task embeds causal information, which leads to superior performance on down stream tasks where capturing this causality matters compared to standard pretraining tasks. As a result, we will not be comparing ourselves with other existing foundation models directly (which are trained with millions of images and heavy gpu resources (Jakubik et al., 2023; Cong et al., 2022)) but rather implement their frameworks (architecture design, pretraining task, etc.) on our dataset to ensure fair comparison and convince researchers in the future to use our approach for pretraining of their remote sensing based foundation models. Note that in our setup the above frameworks can be implemented by changing the inputs and the loss functions, without significant architecture changes. Though MM-MR also uses multi modal data, due to the pretraining task of masked reconstruction, the causal relationship between modalities is not captured in the embeddings and in turn hurts its performance on downstream tasks.
@@ 6.2. Implementation and Pretraining Details
For the pretraining phase, we choose an input series length of 6 images, and selected a random image after the 6th image as the final image to forecast. Since our input series length is 6, we can create multiple samples from one location, i.e one location’s data can result in multiple input series. After splitting the 10000 sampled locations in 60-20-20 split and created numerous samples in each split, We used 50000 image series for training, 10000 images series for validation, and 10000 for testing. We also use 50% spatiotemporally uniform masking for both forecasting and reconstruction based pretraining. We use a patch size of 8 for the vision transformer and a hidden dimension size of 256. Models were trained with a learning rate of 0.0001 on 4 A100 Nvidia GPUs using Adam Optimizer and Mean Squared Error loss. A batch size of 128 was used and all frameworks were trained for 800 epochs each. If the framework required a forecasting phase, it was trained during the final 300 epochs of the total 800 epochs.
@@
All models were trained till convergence. MM-MR reached lower mean squared loss compared to SM-MR, and CI-VSF reached a lower mean square error loss compared to SM-VSF. While the output images produced by SM-MR and MM-MR were quite clean, the images produced by SM-VSF were blurry. CI-VSF was also able to create clean and consistent images, while showing good forecasting performance. For more details please refer to appendix section A.2
7. Results: Soil Moisture Tasks
In this section, we compare the performance of fine-tuned models on soil moisture related tasks, specifically soil moisture forecasting and soil moisture prediction.
7.1. Soil Moisture Forecasting
Dataset and Problem Setting: Soil moisture forecasting plays a important role in the remote sensing domain, enabling various applications such as optimizing crop management, planning irrigation schedules, and improving drought monitoring and mitigation strategies (Dubois et al., 2021; Togneri et al., 2022). In this downstream task, we aim to evaluate the ability of models to forecast soil moisture estimates. Specifically, we seek to predict soil moisture at a given location at an arbitrary future time based on past satellite imagery, weather and soil moisture observations. We focus on 6 Sentinel tiles (T11SKA, T10SEJ, T14SKC, T15TUH, T16SBF, T14RQT) regions (geographic locations shown in Appendix Figure 13) and obtain the corresponding Sentinel2, ERA5 and SMAP (’soil_moisture_am’ band) values for those regions at their optimal temporal and spatial resolutions. Our goal is to forecast SMAP (’soil_moisture_am’ band) values at some point in the future using past context.
To estimate a soil moisture value in the future, the inputs we provide are satellite image time series, weather data till that future date and past soil moisture values. We use a input interval of 6 satellite images and predict the soil moisture value at a 7th timestamp in the future. Each tile is split into 100 grids and a 60-20-20 train-val-test split is done, ensuring no overlap in regions. Multiple samples are generated from each grid, and after preprocessing, we retain 1800 samples for training, 400 for validation, and 400 for testing. To enhance model generalization, we ensure a diverse distribution of forecast range (i.e. the number of days into the future for which forecasts are made), allowing the model to learn both short-term and long-term forecasting patterns. Due to the forecasting related nature of this task, it would be inappropriate to finetune with the SM-MR and MM-MR embeddings, as these methods are not designed for extrapolating into the future. A decoding head for forecasting is appended and finetuned using Mean Absolute Error loss. Encoder is kept fixed. For more architecture and implementation details please refer to Appendix Section A.3.
Table 1. Comparison of models fine-tuned from different pretraining tasks on soil moisture forecasting downstream task. Table reports average Mean Absolute Error (MAE) of forecasted SMAP values for different forecast day ranges. In the test region soil moisture ranged from 0.062 to 0.4998 with a mean of 0.2134 and standard deviation of 0.0997, providing context for interpreting the values below
Soil Moisture Forecasting Downstream Task
Future Forecast Day Range SM-VSF CI-VSF
0 - 25 days 0.0406 0.0179
25 - 50 days 0.0429 0.0184
50 - 100 days 0.0549 0.0189
More than 100 days 0.0678 0.0204
Performance: We train a model for both the SM-VSF and CI-VSF settings using all tiles and a diverse range of forecast day (i.e varying number of days into the future for forecasting). To compare the performance of these models systematically, we compare predictions made across different forecast ranges in the test grids, reported in Table 1. From Table 1, in short forecast ranges (i.e less than 25 days) both SM-VSF and CI-VSF finetuned models are able to forecast soil moisture quite well with a moderate difference between the two. However, when we move to larger forecast ranges, we can see a big jump in the performance of CI-VSF finetuned model compared to SM-VSF’s finetuned model. Even though as the forecast range increases, errors of both models go up, the errors by the SM-VSF model go up significantly higher compared to CI-VSF’s model. This can be attributed to the fact that the CI-VSF model is able to capture the causal relationships between weather satellite imagery and past soil moisture values better than the SM-VSF counterpart, leading to a performance boost especially as we forecast farther into the future.
7.2. Soil Moisture Estimation
Dataset and Problem Setting: Estimation of Soil Moisture at current timestamp using radar imagery has been a topic of interest in the remote sensing community(Adab et al., 2020; Lee et al., 2019; Senanayake et al., 2021). Typically, multiple sources of data including satellite data, radar data, weather data, depth data, etc. are used to estimate in-situ measurements from their site where groundtruth readings are available. In our downstream task, we aim to emulate this task and demonstrate how a pretraining task that captures causal relationships in its embeddings is better for the task of soil moisture estimation. From the same 6 Sentinel tiles used in the soil moisture forecasting problem, we obtain the Sentinel-2 satellite data and ERA5 weather data to serve as our input modalities and the ’soil_moisture_am’ band from SMAP to serve as our groundtruth for soil moisture(Li et al., 2021).
In our problem setting, the input would be a time series of either one or multiple modalities with the goal to estimate the soil moisture at that location for each timestamp. We use a timeframe of 6 satellite images here and aim to estimate the soil moisture at each timestamp. For sample generation, like the previous section, each tile is split into 100 grids and a 60-20-20 train-val-test split is done, and after preprocessing, we retain 1800 samples for training (300 from each tile), 600 for validation (100 from each tile), and 600 for testing (100 from each tile). Due to the prediction related nature of this task, models from all 4 pretraining tasks are finetuned. A decoding head for prediction is appended and finetuned using Mean Absolute Error loss. Encoder is kept fixed. For more architecture and implementation details please refer to Appendix Section A.4.
Performance:
Soil Moisture Prediction Finetuned Models
TEST Tile SM-MR MM-MR SM-VSF CI-VSF
In region Testing: Train on 6 tiles, Test on all 6 tiles
All 0.0615 0.0458 0.0483 0.0282
Cross-Region Testing: Train on 5 tiles, Test on the 6th.
T11SKA 0.1113 0.0847 0.1121 0.0695
T15TUH 0.1283 0.1365 0.1181 0.0834
T14SKC 0.1159 0.1275 0.1312 0.0958
T16SBF 0.0821 0.0631 0.0895 0.0544
T10SEJ 0.1003 0.0718 0.1011 0.0587
T14RQT 0.0815 0.0579 0.0658 0.0558
The performance of finetuned models on this task are presented in Table 2. The first experiment conducted was training using samples from all tiles and testing on samples from the test grids of each tile. From this in region sample testing, we can see CI-VSF’s finetuned model outperforms other models quite significantly. To explore robustness of learned representations, we also explore cross region testing. In this setting, during finetuning, samples from 1 tile are hidden and during testing, models are tested samples from that hidden tile. For example, models are trained on the 1500 samples from 5 tiles (T10SEJ, T14SKC, T15TUH, T16SBF, T14RQT) and tested on the 100 test samples from the hidden tile (T11SKA). We run this setting hiding each tile one at a time, shown in the lower portion of Table 2. From these results, CI-VSF’s model is able to perform cross region soil moisture estimation better than other models in all tiles, showing that embeddings from CI-VSF are able to capture causality and retain this information even when tested on regions not included in finetuning. This highlights how causally informed pretraining strategies help in downstream tasks.

### v4

@@ 4. Pretraining
As mentioned before, our proposed pretraining task is forecasting of satellite imagery using multimodal data in a Knowledge Guided fashion, a pretraining task that varies significantly from the traditional single modality reconstruction. Our proposed approach can be trained using the architecture described above in a direct shot, i.e initialise and update all layers at once. However, this may not be the best way to ensure that the $Emb_{STW}$ embeddings capture causal information, as that is the ultimate aim of our proposed approach. To ensure this information capture, we propose a phase wise pretraining process. Specifically we propose 2 phases of pretraining, where the first phase focuses on getting information from each modality and the second phase focuses on capturing the causal relationship between modalities in the embedding.
@@
To summarize, our proposed pretraining task is to predict a spectral image in the future (response) using a series of spectral images in the past (context) along with the weather till that future date (query). We call this pretraining task as Knowledge Guided Variable step Forecasting (KG-VSF), where the model is expected to forecast k steps into the future. To ensure that relationship across modalities is captured in the embeddings, we adopt a two phase pretraining process. Our hypothesis is that this extra knowledge infusion would help greatly in downstream tasks that rely on land growth and change dynamics such as crop prediction, land cover land use change, etc. A schematic of the stagewise pretraining can be found in appendix A.1. We mask out patches in the spectral imagery and weather data in all the stages to lead to better embeddings, which we describe in a future section(Sec 5.3). An alternative to our two phase pretraining approach would be a single-phase, multi-objective pretraining that simultaneously optimizes masked reconstruction and forecasting losses which might appear more elegant conceptually. However, such a formulation could be suboptimal for our setting where the goal is to embed causal interactions between driver and response modalities rather than just spatiotemporal correlations. For additional discussion, see Appendix A.1.1
5. Dataset Description
@@ 5.1. Data Sources
Our spectral imagery data comes from 6 bands of Sentinel imagery (drusch2012sentinel) and our weather data is from ERA5 land analysis data (hersbach2020era5). We chose these two sources due to their temporal resolutions and more importantly their availability globally from 2021. Our choice of six bands: B2 (Blue), B3 (Green), B4 (Red), B8A (Narrow NIR), B11 (SWIR-1), and B12 (SWIR-2) was made for both scientific and practical reasons. Prior studies (e.g.,(jakubik2023foundationmodelsgeneralistgeospatial), (brown2025alphaearthfoundationsembeddingfield)) have also used these exact bands to demonstrate good performance on a number of downstream tasks ranging from agricultural to climate oriented tasks. Including all 13 bands would more than double the amount of storage space required. Working with fewer bands improved model training speed, and decreased model architecture size, both of which were critical given our limited computational resources
For each satellite image instance, we also collect the day of the year it came from, thus forming a series with values from 1 to 365 and a length the same as the number of image instances for that location. Though ERA5 data source consists of various bands that are useful for land cover related tasks we chose 5 bands that are the primary weather drivers, namely (temperature 2m min, temperature 2m max, total precipitation sum, $u$ component of wind 10m, $v$ component of wind 10m). ERA5 data is available at a daily temporal resolution and a spatial resolution of  11k meters.
To summarize, for each location our data sources are:
$\bullet$ Spectral Imagery Series: A series of Sentinel2 Imagery of 6 bands. Length of this series depends on coverage of the location.
$\bullet$ Weather Data Series: A series of ERA5 Land data of 5 bands , with a series length of 365 (one per day).
@@
5.2. Datasets
We constructed two datasets: a Global Dataset and a US Cropland Dataset to serve different objectives. The Global Dataset includes samples from diverse regions across the world, providing wide spectral variability that benefits tasks requiring generalizable spectral representations. In contrast, the US Cropland Dataset focuses on agriculturally active regions within the United States, enabling the model to better learn patterns specific to crop growth and phenological changes. This division allows us to examine how pretraining on broad versus domain-specific data influences the quality and transferability of learned embeddings across different downstream applications. Below we describe the sampling strategy and distribution for both these datasets:
Global Dataset: We begin by randomly selecting 10,000 unique land locations across the globe, each covering a 128×128 Sentinel-2 (approximately 1km x 1km area). These 10,000 locations are divided into training, validation, and testing splits in a 60:20:20 ratio, corresponding to 6,000, 2,000, and 2,000 locations, respectively. For each location, all available Sentinel-2 images for the year 2021 are collected—typically ranging from 40 to 70 timestamps depending on regional coverage (e.g., 70+ images in US and Europe and about 40 images in India). For the pretraining phase, we choose an input series length of 6 images, and select a random image after the 6th image as the image to forecast. This requires sequences of 7 timestamps for each sample that can be used for training, validation, or testing. Since a single location can theoretically produce multiple unique 7-timestamp sequences (e.g., for a location with 50 timestamps, one could create 50-choose-7, i.e approximately 99.8 million possible combinations), we subsample a small number of representative temporal sequences per location to maintain diversity while keeping the dataset size tractable. Specifically, for each of the 6,000 training locations, we randomly sample 7-timestamp sequences 10 times, yielding a total of 60,000 possible samples from which 50,000 are randomly chosen for pretraining. Similarly, for each of the 2,000 validation and 2,000 testing locations, we sample 5 7-timestamp sequences per location, resulting in 10,000 samples in each split. This results in an overall 50k/10k/10k split. This approach ensures that while spatial coverage remains balanced across splits, the model is also exposed to rich temporal variability within each region.
US Cropland Dataset: We begin by selecting and downloading 50 Sentinel tiles across the contiguous US for the year of 2021, focusing on tiles that contain a high amount of crop cover. Each Sentinel tile covers a region of 10980x10980 pixels at a 10m resolution (approximately 110km x 110km area). Since our sample analysis size is 128x128, numerous non overlapping 128x128 regions can be selected from the overall 10980x10980 region, specifically approximately 85x85 (=7225) non overlapping 128x128 regions can be created. Also, similar to the Global Dataset, for each tile, all available Sentinel-2 images for the year 2021 are collected and filtered to ensure that their cloud cover/missing values are less than 20%. Similar to the Global data set, we sample 7-timestamp sequences for each 128x128 region for each of the sentinel tiles. Specifically, we randomly select a 128x128 region from each sentinel tile, and then further identify a unique 7-timestamp sequence for that region to create a sample. Selection of 128x128 regions are biased towards the regions that have more crop coverage. For each tile in our dataset, we create 2000 such samples, resulting in a total of 100000 samples(50 tiles* 2000 per tile). We then choose 30 tiles for training and 10 tiles for validation, and 10 tiles for testing, resulting in a 60k/20/20k data split. This approach ensures that there is spatial coverage across different cropland areas in the US.
Since each sample from both datasets covers an area of 1.28km x 1.28km(i.e 128 x 128 Sentinel Pixels), The weather data per timestamp will be of shape 1x1, i.e one value per weather band per day. The reason behind having only one pixel value from weather data is due to its coarse resolution (11km) compared to the area coverd by our Sentinel patch (1.28km x 1.28km)
@@ 5.3. Masking
Since we have a spatiotemporal input and architecture, we adopt a spatiotemporally uniform masking method, i.e, masking that is fair both spatially and temporally. In our unique masking strategy, there are an equal number of masked patches per timestamp as well as an equal number of masked patches per spatial patch location along the temporal axis. This ensures that all temporal patch series that can be created from unmasked patches at each spatial location would be of same length, easing the implementation of temporal components. Also the shared vision transformer will also have same number of outputs per timestamp due to same number of unmasked patches per timestamp. For a visual representation of this masking, please refer to the Appendix Section A.2.
@@ 6.1. Comparative Pretraining Frameworks
In our experiments, we evaluate the effectiveness of KG-VSF by comparing with several representative foundation modeling frameworks. We model these frameworks by varying the choices for input and pretraining tasks from existing works, as described below.
$\bullet$ SM-MR: Single Modality Masked Reconstruction (SM-MR) with satellite data as input and MR as the pretraining task. This is the most common setup used in existing remote sensing foundation models (e.g.,(jakubik2023foundationmodelsgeneralistgeospatial; cong2022satmae; sun2022ringmo)).
$\bullet$ MM-MR: MultiModal Masked Reconstruction (MM-MR) with satellite and weather data as input series and MR as pretraining task. This represents the foundation models that use multiple modalities but still do masked reconstruction (eg. (nedungadi2024mmearth; tseng2023lightweight))
$\bullet$ SM-VSF: Single Modality Variable step Forecasting (SM-VSF) with the satellite data as input series and Variable Step Forecasting(VSF) as pretraining task.
$\bullet$ MM-VSF: Multi Modality Variable step Forecasting (MM-VSF) with the satellite data and weather data as input series and Variable Step Forecasting(VSF) as pretraining task.
It is important to note that the goal of the paper is not to release a new foundation model but show efficacy of our novel pre-training strategy for building foundation models. We are aiming to show how a Knowledge Guided pretraining task embeds causal knowledge information, which leads to superior performance on downstream tasks where capturing this causality matters compared to standard pretraining tasks. As a result, we will not be focusing on comparing ourselves with other existing foundation models directly (which are trained with millions of images and heavy gpu resources (jakubik2023foundationmodelsgeneralistgeospatial; cong2022satmae)) but rather implement their frameworks (architecture design, pretraining task, etc.) on our dataset to ensure fair comparison and convince researchers in the future to use our approach for pretraining of their remote sensing based foundation models. For some of the downstream tasks, we do include task specific models and also comparison with Prithvi EO.2 foundation model (szwarcman2025prithvieo20versatilemultitemporalfoundation). Note that in our setup the above frameworks can be implemented by changing the inputs and the loss functions, without significant architecture changes. Though MM-MR also uses multi modal data, due to the pretraining task of masked reconstruction, the causal knowledge relationship between modalities is not captured in the embeddings and in turn hurts its performance on downstream tasks. We also expect a similar scenario with the MM-VSF pretraining framework as well.
@@ 6.2. Implementation and Pretraining Details
As mentioned in the Dataset Section 5.2, For the pretraining phase, we choose an input series length of 6 images, and selected a random image after the 6th image as the final image to forecast. We also use 50% spatiotemporally uniform masking for both forecasting and reconstruction based pretraining(For reasoning behind this choice of 50% please refer to Appendix Section A.2). We use a patch size of 8 for the vision transformer and a hidden dimension size of 256. This means for each 8x8 patch we would get an embedding of size 256. Models were trained with a learning rate of 0.0001 on 4 A100 Nvidia GPUs using Adam Optimizer and Mean Squared Error loss. A batch size of 128 was used and all frameworks were trained for 800 epochs each. If the framework required a forecasting phase, it was trained during the final 300 epochs of the total 800 epochs.
@@
All models were trained till convergence. MM-MR reached lower mean squared loss compared to SM-MR, and KG-VSF reached a lower mean square error loss compared to SM-VSF and MM-VSF. While the output images produced by SM-MR and MM-MR were quite clean, the images produced by SM-VSF and MM-VSF were blurry. KG-VSF was also able to create clean and consistent images, while showing good forecasting performance. For more details please refer to appendix section A.3.
For each downstream task, only the final layers are fine-tuned, while the layers used to create the embeddings are kept fixed. This design allows a fair comparison of different pre-training objectives in terms of their ability to produce informative embeddings. Consequently, the performance observed on the downstream tasks reflects how effectively each pretext task captures relevant information from the input modalities during the embedding process. For the global data set, the chances of overlap between data used in any of our downstream tasks and 50,000 locations globally is very slim. For the US Cropland data set, all the data sets used in downstream tasks only come from the 10 test tiles that were never used for pre-training (and validating) the models for creating embeddings.